2-sinQ cosQ- cosQ /1-sinQ+ sin2Q-cos2Q =cotQ solve by trigonometrical identities tignometry class 10

Grade:10

1 Answers

Avinash
5 years ago
Cross multiply;2-sinQ.cosQ-cosQ=cotQ-cosQ+2cos^Q-cos2Q.cotQ==2-sinQ.cosQ=cotQ[1-cos2Q]+2cos^Q=cotQ [2sin^Q] +2cos^Q==2-sinQcosQ=2sinQcosQ+2Cos^Q==2-2cos^Q=3sinQcosQ==2[sin^Q]=3sinQcosQ==2sinQ=3cosQtanQ=3/2Therefore sinQ=3/√13CosQ=2/√13